The plates of a parallel plate capacitor,5.0x10^-3m apart are maintained at a potential difference of 5.0x10^4 Calculate the magnitude of the I. Electric field intensity between the plates II. Force on the electron.
1
Expert's answer
2020-05-12T09:39:02-0400

Given, the separation between parallel plate of capacitor is "d=5.0 \times 10^{-3} \: m" , the potential difference is "V=5.0 \times 10^4" volt.

Now , we know that

"V=E\cdot d \\\implies E=\frac{V}{d}"

Where, "E" denotes the electric field intensity inside the capacitor.

i). Now, from the above formula we get,


"E=\frac{5.0 \times 10^4}{5.0 \times 10^{-3}} \: \text{Newton/Columb}(N/C)\\
\implies E=1.0 \times10^7 \: N/C"

Therefore, Electric field intensity inside the capacitor is "E=1.0 \times10^7 \: N/C" .



ii). Since, force acting on a charge due to electric field is given by

"F=e \cdot E"

and S.I unit is Newton "(N)", where "e = 1.602 \times 10^{-19} C" is the magnitude of the charge on a electron.

Hence, force acting on the electron is,

"F=1.602 \times 10^{-19} \cdot( 1.0 \times 10^7)\\
\implies F=1.602\times10^{-12} \: N"

in the opposite direction of electric field.

Hence, we are done.